OpenStudy (anonymous):

Nine less than 4 times a number is the same as twice the number increased by 11. What is the unknown number?

OpenStudy (ribhu):

4x-9 = 2x+11

OpenStudy (hockeychick23):

4x-9= 2x+11 Lets add 9 to both sides 4x-9+9= 2x+11+9 4x= 2x+20 Subtract 2x from both sides 4x-2x=2x-2x+20 2x=20 x=10

OpenStudy (ribhu):

where x would be the number, this would be the linear equation required.

OpenStudy (ribhu):

let them solve on their own!!

OpenStudy (anonymous):

WHAT

OpenStudy (hockeychick23):

Ok so your equation says: nine less than 4 times a number is the same as twice the number increased by 11. What is the unknown number? so we are using x as "the number" so the first part of the equation says 9 less than 4 times the number so we do 4x-9, do you get that so far?

OpenStudy (anonymous):

no not really

OpenStudy (hockeychick23):

ok so do you get that we used x as our variable?

OpenStudy (anonymous):

no just tell me the answer

OpenStudy (hockeychick23):

i already did i said x=10

OpenStudy (anonymous):

oh my bad sorry and thnx

OpenStudy (hockeychick23):

no problem